# Inductive Domain Transfer In Misspecified Simulation-Based Inference

### v2

@@ 2 Background
2.1 Simulation-Based Inference and Neural Posterior Estimation
@@
Very often, when ${\bm{x}}_{s}$ is high-dimensional, a neural statistics encoder (NSE, wehenkel2024addressing ) $h_{\omega}$ is used to obtain a lower-dimensional representation with sufficient information for the posterior inference task of parameters $\bm{\theta}$. Eventually, in neural posterior estimation (NPE), one minimizes the expected negative log-likelihood over simulator draws w.r.t parameters $\bm{\theta}$ :
@@
2.2 Transductive Semi-Supervised Posterior Estimation (RoPE)
@@
Once the NSE, $h_{\omega^{\star}}$, and NPE, $q_{\psi^{\star}}$ from Section 2.1 are trained and fixed, a small set (calibration set) of real observations ${\bm{x}}_{r}^{i}$ and their corresponding known parameters $\bm{\theta}_{i}$, is used to adapt the encoder to the domain shift. Each $\bm{\theta}_{i}$ is passed through the simulator to obtain the corresponding ${\bm{x}}_{s}^{i}=S(\bm{\theta}_{i})$ and the calibration set becomes $\mathcal{D}_{calib}=\{{\bm{x}}_{r}^{i},{\bm{x}}_{s}^{i}\}_{i=1}^{N_{calib}}$, where $N_{calib}$ is the number of samples in the set. Then, $g_{\phi}$, the NSE for the real observations, initialized as $h_{\omega^{\star}}$, is fine-tuned on $\mathcal{D}_{calib}$ to minimize the mean squared error between the embedding of a real observation, $g_{\phi}({\bm{x}}_{r}^{i})$, and the embedding of its corresponding simulated sample $h_{\omega^{\star}}({\bm{x}}_{s}^{i})$, for every pair of ${\bm{x}}_{r}^{i},{\bm{x}}_{s}^{i}\in\mathcal{D}_{calib}$. The outcome is a limited (depending on the size of $\mathcal{D}_{calib}$) domain adaptation of the NSE $g_{\phi}$, which permits to encode representations of real observations in the same latent space as $h_{\omega^{\star}}(\mathbf{x}_{s})$.
@@
Since NSE fine-tuning relies on a limited calibration set, its ability to generalize to unseen real observations is constrained, leaving residual uncertainty in the domain transfer. RoPE wehenkel2024addressing takes this uncertainty into account by coupling real and simulated observations through entropic OT, which provides soft matching between embeddings $\{g_{\phi^{\star}}({\bm{x}}_{r}^{i})\}$ of $N_{test}$ test observations $\mathcal{D}_{test}=\{{\bm{x}}_{r}^{i}\}_{i=1}^{N_{test}}$ and embeddings $\{h_{\omega^{\star}}({\bm{x}}_{s}^{j})\}$ of a fresh simulation set $\mathcal{D}_{OT}=\{{\bm{x}}_{s}^{j}\}_{j=1}^{N_{OT}}$ with $N_{OT}$ samples. Specifically, it solves the following semi-balanced entropic OT problem:
@@
where $\bm{P}$}, the transport plan, is constrained row-wise to $\mathcal{B}(N_{test},N_{OT})=\{{\bm{P}}\in\mathbb{R}_{+}^{N_{test}\times N_{OT}}|{\bm{P}}\bm{1}_{N_{OT}}=\frac{\bm{1}}{N_{test}}\bm{1}_{N_{test}}\}$ and $\bm{C}$} is the pairwise euclidean distance matrix between both sets of embeddings. The weight $\rho$ governs the relaxation of the column-marginal constraint via a Kullback–Leibler divergence (KL) penalty, enabling unbalanced OT to accommodate prior misspecification, while $\gamma$ controls the entropy regularization strength, thereby tuning the method’s sensitivity to model misspecification and to uncertainty introduced during encoder fine-tuning on the calibration set. Problem 1 is commonly solved using an iterative Bregman projection solver benamou2015iterative ; frogner2015learning ; vincentsemi ; sejourne2023unbalanced . As detailed in the Appendix, it comes down to update along iterations $t$ the transport plan $\bm{P}$}, following:
@@
Finally, the calibrated posterior for each ${\bm{x}}_{r}^{i}$ is
@@ 3 Methods
We propose a new framework for robust inductive domain transfer in misspecified simulation-based inference (FRISBI). It relies on a unified workflow that achieves a joint distribution-level and point-wise alignment while enabling inductive inference, thereby extending the solution to misspecified SBI beyond the limitations described above. The encoder $g_{\phi}$ is first trained using a joint objective that combines a variant of entropic OT admitting closed-form solutions, with a supervised calibration loss, as detailed in Section 3.1. Then to avoid the need for accessing simulations at test time, we further amortize this solution using the inductive strategy presented in Section 3.2. A complete description of the full pipeline and training procedure is provided in Algorithm 3.2.
@@ 3.1 Balancing Unpaired Alignment and Point-wise Domain Transfer
We specifically enforce $\rho=0$ in the entropic OT objective as the resulting optimization problem w.r.t $\bm{P}$} is naturally well-suited for inductive learning. Indeed, one can see that setting $\rho=0$ in Eq. (1), implies that this problem admits a closed-form solution ${\bm{P}}^{\star}=\text{diag}(\frac{\bm{1}_{N_{u}}}{N_{u}{\bm{K}}\bm{1}}){\bm{K}}$ where ${\bm{K}}=e^{-{\bm{C}}/\gamma}$ and $\bm{C}$} is the pairwise euclidean distance matrix between embeddings (see also (benamou2015iterative, , Proposition 1)). This leads to the efficient stochastic gradient descent (SGD) algorithm described in Stage 1 of Algorithm 3.2, which alternates between computing embeddings for real-world observations and computing these closed-form coupling independently for each embedding. Remark that an analog strategy could be performed when $\rho>0$, replacing the closed-form computation with numerous computations of the update rule in Eq. 2 until convergence, as often done in mini-batch OT fatras2021minibatch ; fatras2021unbalanced . While being more computationally intensive for training, such approaches are also known for leading to biased OT solutions and a significant sensitivity to the chosen batch size. Hence, it further motivates our choice for setting $\rho=0$ in Eq. 1 in our mini-batch OT formulation.
@@ 4.1 Benchmarks
This benchmark uses a subset sun2005estimating ; goldberger2000physiobank of the MIMIC-II dataset saeed2011multiparameter , comprising 350 patients who underwent thermodilution—a procedure estimating cardiac output (CO) via cold fluid injection and downstream temperature measurement. Each patient has arterial blood pressure (ABP) signals aligned with CO readings, yielding $\sim 2200$ valid ABP segments with corresponding CO values. For simulation, we use OpenBF Benemerito_2024 , a validated 1D cardiovascular flow simulator supporting fast, multiscale finite-volume simulations. The estimated parameters are $\bm{\theta}=\{\text{HR},\text{CO}\}$, where HR is heart rate. The empirical means and standard deviations of HR and CO in the dataset are $(87,12)$ beats/min and $(5.1,1.6)$ L/min respectively. In contrast, the CO from the simulations is derived by the known connection $\text{CO}=\text{HR}\times\text{SV}$, where SV is the stroke volume. HR and SV are sampled from uniform distributions $\text{U}(50,150)$ beats/min and $\text{U}(40,140)$ L/beat respectively. This yields CO in the range of $[2,20]$ L/min.
@@ 4.2 Experimental design
The primary method we compare against is RoPE wehenkel2024addressing . However, a direct comparison on the same test set is unfair, as RoPE is purely transductive. Nevertheless, we include this setting as a baseline, denoted RoPE full test. For a fairer inductive comparison, we introduce a single-sample variant of RoPE: for each test point ${\bm{x}}_{r}^{\text{test}}$, we add it to the unpaired training set $\mathcal{D}_{u}$, compute the OT coupling with simulations $\mathcal{D}_{OT}$, and estimate the posterior using the resulting plan. This is repeated per test point, and we denote this baseline as RoPE single sample. We also compare against baselines from wehenkel2024addressing , including NPE (see Section 2.1), applied directly to real observations without domain transfer. In addition, we include OT-only baselines that do not adapt the embedding space. Here, the fixed encoder $h_{\omega^{\star}}$ is applied to both real and simulated samples, and OT is computed in this space. The full-test variant is denoted OT-only (full test), and the single-sample variant as OT-only (single sample). For all experiments involving OT, we use solvers from the POT Python library flamary2021pot ; flamary2024pot .
@@
The second metric is the average coverage area under the curve (ACAUC), which reflects how well the model’s credible intervals align with the true parameter coverage. It provides a measure of calibration by comparing the fraction of true parameters falling within the estimated credible intervals at different confidence levels. A perfectly calibrated model has an ACAUC of zero, while positive values indicate overconfident estimates and negative values indicate underconfident estimates. For a detailed mathematical definition, we refer the reader to wehenkel2024addressing .
@@ 4.3.1 Performance Across Different Calibration Set Sizes
The results, including mean scores and standard deviations computed across folds, are presented in Fig. 2. Our method (solid red line) consistently outperforms the transductive single-sample RoPE baseline (solid orange line) in terms of LPP, while maintaining ACAUC close to zero across most calibration set sizes. This indicates that it achieves a robust inductive parameter estimation compared to RoPE. With larger calibration sets, our method matches or even exceeds the performance of the full-test RoPE (solid green line), which has access to the entire test set—highlighting the effectiveness of combining OT-based domain transfer with supervised calibration. However, as the calibration set size increases, our method—like the fine-tuning baselines—shows a decline in confidence (i.e., lower ACAUC), suggesting that the joint training procedure increasingly relies on the supervised loss over the OT loss as more calibration data becomes available. In contrast, the transductive RoPE baseline is less affected by this trend, possibly because it is applied directly to the test set. Nevertheless, compared to the single-sample version of RoPE, which consistently exhibits overconfidence, our method achieves noticeably better calibration.
@@ 5 Conclusion
In this chapter, we discuss model-misspecification correction through two complementary approaches to physics-informed latent parameter inference: hybrid ML and SBI. While hybrid ML appears to be the more straightforward choice for correcting misspecification—since we simply introduce a data-driven correction model into the forward model (i.e., the likelihood $p(\mathbf{x}\mid\bm{\theta})$)—we observed that, despite promising preliminary results in inferring latent cardiac biomarkers such as heart rate (HR) and left ventricular ejection time (LVET), this approach suffered from mode collapse when applied to another cardiac marker, cardiac output (CO). The collapse arose from identifiability issues in the surrogate model. This suggests that hybrid-ML-based corrections are most effective when (1) the simulation can be explicitly embedded within the hybrid model or (2) the surrogate accurately approximates the true simulation.
@@
To address these limitations, we shift focus to SBI, which avoids the need to amortize or explicitly model the forward process (i.e., the likelihood), thereby removing reliance on a surrogate model. The challenge then becomes designing a principled method to correct for the domain shift between real-world and simulated observations. To this end, we introduce an amortized and inductive domain-transfer framework for posterior estimation that accounts for misspecification in simulation-based inference. Our method leverages mini-batch optimal transport to enable joint training on both unpaired data and small paired calibration sets. The final posterior, approximated via an optimal transport-based mixture, is amortized using a conditional normalizing flow. This eliminates the need for additional transport computations or access to simulations at test time—a key limitation of transductive approaches such as RoPE. Our approach demonstrates competitive performance across a range of benchmarks, including both synthetic and complex real-world datasets, and specifically in the inference of CO within a real clinical cohort.
@@
Limitations and future directions. Sensitivity to label noise, as observed in our experiments, suggests that the quality of training data significantly impacts posterior accuracy. While increasing the entropy regularization weight may offer partial mitigation, more adaptive loss formulations—such as dynamically adjusting the supervised loss weight $\lambda$—should be explored. Additionally, active learning or uncertainty-aware sampling strategies could be employed to guide calibration set selection under a fixed annotation budget.
Furthermore, our training procedure currently consists of two separate stages: (i) joint optimal transport and supervised training, and (ii) amortization of the induced posterior mixture. A more unified alternative could involve learning the transport mapping directly through a neural optimal transport framework, potentially integrating both matching and inference into a single end-to-end process.
Appendix A Semi-balanced Optimal transport solvers.

### v3

@@ 2 Background
2.1 Simulation‐Based Inference and Neural Posterior Estimation
@@
Very often, when ${\bm{x}}_{s}$ is high-dimensional, a neural statistics encoder (NSE, wehenkel2024addressing ) $h_{\omega}$ is used to obtain a lower-dimensional representation with sufficient information for the posterior inference task of parameters $\bm{\theta}$. Eventually, in neural posterior estimation (NPE), one minimizes the expected negative log‐likelihood over simulator draws w.r.t parameters $\bm{\theta}$ :
@@
2.2 Transductive Semi‐Supervised Posterior Estimation (RoPE)
@@
Once the NSE, $h_{\omega^{\star}}$, and NPE, $q_{\psi^{\star}}$ from Section 2.1 are trained and fixed, a small set (calibration set) of real observations ${\bm{x}}_{r}^{i}$ and their corresponding known parameters $\bm{\theta}_{i}$, is used to adapt the encoder to the domain shift. Each $\bm{\theta}_{i}$ is passed through the simulator to obtain the corresponding ${\bm{x}}_{s}^{i}=S(\bm{\theta}_{i})$ and the calibration set becomes $\mathcal{D}_{calib}=\{{\bm{x}}_{r}^{i},{\bm{x}}_{s}^{i}\}_{i=1}^{N_{calib}}$, where $N_{calib}$ is the number of samples in the set. Then, $g_{\phi}$, the NSE for the real observations, initialized as $h_{\omega^{\star}}$, is fine‐tuned on $\mathcal{D}_{calib}$ to minimize the mean squared error between the embedding of a real observation, $g_{\phi}({\bm{x}}_{r}^{i})$, and the embedding of its corresponding simulated sample $h_{\omega^{\star}}({\bm{x}}_{s}^{i})$, for every pair of ${\bm{x}}_{r}^{i},{\bm{x}}_{s}^{i}\in\mathcal{D}_{calib}$. The outcome is a limited (depending on the size of $\mathcal{D}_{calib}$) domain adaptation of the NSE $g_{\phi}$, which permits to encode representations of real observations in the same latent space as $h_{\omega^{\star}}(\mathbf{x}_{s})$.
@@
Since NSE fine-tuning relies on a limited calibration set, its ability to generalize to unseen real observations is constrained, leaving residual uncertainty in the domain transfer. RoPE takes this uncertainty into account by coupling real and simulated observations through entropic OT, which computes a soft assignment matrix between embeddings $\{g_{\phi^{\star}}({\bm{x}}_{r}^{i})\}$ of test observations $\mathcal{D}_{test}=\{{\bm{x}}_{r}^{i}\}_{i=1}^{N_{test}}$ and embeddings $\{h_{\omega^{\star}}({\bm{x}}_{s}^{j})\}$ of a fresh simulation set $\mathcal{D}_{OT}=\{{\bm{x}}_{s}^{j}\}_{j=1}^{N_{OT}}$. The latter then acts as prototypes to which the matched embeddings $\{g_{\phi^{\star}}({\bm{x}}_{r}^{i})\}$ must be close in an Euclidean sense, similarly to the soft Kmeans algorithm canas2012learning . Specifically, they propose to solve for the following semi-balanced entropic OT problem vincentsemi ; van2024graph :
@@
where $\bm{P}$}, the transport plan, is constrained row‐wise to $\mathcal{B}(N_{test},N_{OT})=\{{\bm{P}}\in\mathbb{R}_{+}^{N_{test}\times N_{OT}}|{\bm{P}}\bm{1}_{N_{OT}}=\frac{\bm{1}}{N_{test}}\bm{1}_{N_{test}}\}$ and $\bm{C}$} is the pairwise euclidean distance matrix between both sets of embeddings. The weight $\rho$ encourages prototypes to be matched to a uniform number of test samples, enabling unbalanced OT to accommodate prior misspecification, while $\gamma$ controls the entropy regularization strength, thereby tuning the method’s sensitivity to model misspecification and to uncertainty introduced during encoder fine‐tuning on the calibration set. Problem 1 is commonly solved using an iterative Bregman projection solver benamou2015iterative ; frogner2015learning ; vincentsemi ; sejourne2023unbalanced . As detailed in the Appendix, it comes down to update along iterations $t$ the transport plan $\bm{P}$}, following:
@@
Finally, wehenkel2024addressing shows that the calibrated posterior for each ${\bm{x}}_{r}^{i}$ can be approximated by marginalizing over ${\bm{x}}_{s}$, resulting in the following posterior:
@@ 3 Methods
We propose a new Framework for Robust Inductive domain transfer in misspecified Simulation-Based Inference (FRISBI). It relies on a unified workflow that achieves a joint distribution-level and point-wise alignment while enabling inductive inference, thereby extending the solution to misspecified SBI beyond the limitations described above. The encoder $g_{\phi}$ is first trained using a joint objective that combines a variant of entropic OT admitting closed-form solutions, with a supervised calibration loss, as detailed in Section 3.1. Then to avoid the need for accessing simulations at test time, we further amortize this solution using the inductive strategy presented in Section 3.2. A complete description of the full pipeline and training procedure is provided in Algorithm 3.2. For clarity, Appendix A includes a summary table describing the different datasets.
@@ 3.1 Balancing Unpaired Alignment and Point-wise Domain Transfer
We specifically enforce $\rho=0$ in the entropic OT objective as the resulting optimization problem w.r.t $\bm{P}$} is naturally well-suited for inductive learning. Indeed, one can see that setting $\rho=0$ in Eq. (1), implies that this problem admits a closed-form solution ${\bm{P}}^{\star}=\text{diag}(\frac{\bm{1}_{N_{u}}}{N_{u}{\bm{K}}\bm{1}}){\bm{K}}$ where ${\bm{K}}=e^{-{\bm{C}}/\gamma}$ and $\bm{C}$} is the pairwise euclidean distance matrix between embeddings (see also (benamou2015iterative, , Proposition 1)). This leads to the efficient stochastic gradient descent (SGD) algorithm described in Stage 1 of Algorithm 3.2, which alternates between computing embeddings for real-world observations and computing these closed-form coupling independently for each embedding. Setting $\rho=0$ enables this closed-form computation, requiring $N_{\text{OT}}$ operations per sample in $D_{u}$, each involving a Euclidean distance in $\mathbb{R}^{d}$, for an overall complexity of $\mathcal{O}(d)$. In the mini-batch setting, this yields a per-step complexity of $B_{t}N_{\text{OT}}d$, where $B_{t}$ is the batch size. Remark that an analogous strategy could be applied when $\rho>0$, replacing the closed-form computation with iterative updates of Eq. 2 until convergence, as typically done in mini-batch OT fatras2021minibatch ; fatras2021unbalanced . However, this approach is more computationally demanding and prone to bias, with high sensitivity to batch size. In addition, RoPE, which uses $\rho>0$ and cannot operate on mini-batches, scales as $\mathcal{O}(\log(N_{u}N_{\text{OT}})N_{u}N_{\text{OT}}d)$, where $N_{u}$ is the total number of real observations. These considerations further motivate our choice of $\rho=0$ for improved efficiency and scalability.
@@ 4.1 Benchmarks
This benchmark uses a subset sun2005estimating ; goldberger2000physiobank of the MIMIC-II dataset saeed2011multiparameter , comprising 350 patients who underwent thermodilution—a procedure estimating cardiac output (CO) via cold fluid injection and downstream temperature measurement. Each patient has arterial blood pressure (ABP) signals aligned with CO readings, yielding $\sim 2200$ valid ABP segments with corresponding CO values. For simulation, we use OpenBF Benemerito_2024 , a validated 1D cardiovascular flow simulator supporting fast, multiscale finite-volume simulations. The estimated parameters are $\bm{\theta}=\{\text{HR},\text{CO}\}$, where HR is heart rate. The empirical means and standard deviations of HR and CO in the dataset are $(87,12)$ beats/min and $(5.1,1.6)$ L/min respectively. In contrast, the CO from the simulations is derived by the known connection $\text{CO}=\text{HR}\times\text{SV}$, where SV is the stroke volume. HR and SV are sampled from uniform distributions $\text{U}(50,150)$ beats/min and $\text{U}(40,140)$ L/beat respectively. This yields CO in the range of $[2,20]$ L/min. This use case is inherently affected by label noise, since the pairing of HR and CO values with observed pulse waves depends on temporal alignment through patients’ electronic records, which is prone to misalignments.
@@ 4.2 Experimental design
The primary method we compare against is RoPE wehenkel2024addressing . However, a direct comparison on the same test set is unfair, as RoPE is purely transductive. Nevertheless, we include this setting as a baseline, denoted RoPE full test. For a fairer inductive comparison, we introduce a single-sample variant of RoPE: for each test point ${\bm{x}}_{r}^{\text{test}}$, we add it to the unpaired training set $\mathcal{D}_{u}$, compute the OT coupling with simulations $\mathcal{D}_{OT}$, and estimate the posterior using the resulting plan. This is repeated per test point, and we denote this baseline as RoPE single sample. We also compare against baselines from wehenkel2024addressing , including NPE (see Section 2.1), applied directly to real observations without domain transfer. To assess the role of the calibration set, we also include an unsupervised domain adaptation (UDA) NPE baseline following swierc2024domain . In addition, we include OT-only baselines that do not adapt the embedding space. Here, the fixed encoder $h_{\omega^{\star}}$ is applied to both real and simulated samples, and OT is computed in this space. The full-test variant is denoted OT-only (full test), and the single-sample variant as OT-only (single sample). For all experiments involving OT, we use solvers from the POT Python library flamary2021pot ; flamary2024pot .
@@
The second metric is the average coverage area under the curve (ACAUC), which reflects how well the model’s credible intervals align with the true parameter coverage. It provides a measure of calibration by comparing the fraction of true parameters falling within the estimated credible intervals at different confidence levels. In all experiments, credible intervals are obtained by drawing $1000$ samples from the corresponding approximate posterior distribution. A perfectly calibrated model has an ACAUC of zero, while positive values indicate overconfident estimates and negative values indicate underconfident estimates. For a detailed mathematical definition, we refer the reader to wehenkel2024addressing .
@@ 4.3.1 Performance Across Different Calibration Set Sizes
The results, including mean scores and standard deviations computed across folds, are presented in Fig. 2. Our method (solid red line) consistently outperforms the transductive single-sample RoPE baseline (solid orange line) in terms of LPP, while maintaining ACAUC close to zero across most calibration set sizes. This indicates that it achieves a robust inductive parameter estimation compared to RoPE. With larger calibration sets, our method matches or even exceeds the performance of the full-test RoPE (solid green line), which has access to the entire test set—highlighting the effectiveness of combining OT-based domain transfer with supervised calibration. It also highlights the importance of incorporating a calibration set, even a small one, as the UDA baseline completely fails under significant misspecification, consistent with the observations in elsemuller2025does . However, as the calibration set size increases, our method—like the fine-tuning baselines—shows a decline in confidence (i.e., lower ACAUC), suggesting that the joint training procedure increasingly relies on the supervised loss over the OT loss as more calibration data becomes available. In contrast, the transductive RoPE baseline is less affected by this trend, possibly because it is applied directly to the test set. Nevertheless, compared to the single-sample version of RoPE, which consistently exhibits overconfidence, our method achieves noticeably better calibration. We also evaluate performance on two additional benchmarks exhibiting more moderate misspecification (Appendix C.1), also considered in wehenkel2024addressing . Similar to RoPE, FRISBI shows no significant advantage over baseline methods in simpler and minimally misspecified setting.
@@ 4.3.2 Ablation
Hyperparameter sensitivity analysis. In Appendix C.2, we present a sensitivity analysis of the hyperparameters $\gamma$ and $\lambda$ on the Light Tunnel benchmark. As in RoPE, a larger $\gamma$ is beneficial when the calibration set is small, producing a more diffuse OT coupling. With larger calibration sets (e.g., 1,000 samples), $\gamma$ can be reduced to obtain a sharper coupling. The effect of $\lambda$ is evaluated under 10% label noise, since noise influences the weighting of the supervised objective. Higher $\lambda$ values decrease confidence as the calibration set grows, while smaller values yield better LPP scores, suggesting that an adaptive tuning of $\lambda$ based on noise level and data size may be advantageous.
@@ 5 Conclusion
In this work, we introduce an amortized and inductive posterior estimator for misspecified simulation-based inference. Our method leverages mini-batch optimal transport to enable joint training over both unpaired and small paired calibration sets. The final posterior, approximated by an OT-based mixture, is amortised by a conditional normalizing flow, eliminating the need for additional transport computations or access to simulations at test time – a key limitation of transductive approaches like RoPE. Our approach demonstrates competitive performance across a range of benchmarks, including both synthetic and complex real-world datasets.
Dimensionality and scalability. While our experiments focus on relatively low-dimensional cases representative of many real-world scenarios, our framework naturally extends to higher-dimensional settings. Two main challenges arise: (i) the mismatch between simulated and real embeddings tends to increase with dimensionality, requiring larger calibration sets, and (ii) the embedding dimension needed to capture sufficient statistics typically grows with the parameter dimension, complicating the coupling. Following Chen et al. chen2020neural , we use an overparameterized embedding dimension of 16 to mitigate this issue. The joint optimization of supervised and OT objectives further alleviates the limitations of small calibration sets by exploiting unlabeled data. As discussed in Section 3.1, FRISBI is expected to scale more efficiently than RoPE in high-dimensional settings.
@@
Limitations and future directions. Sensitivity to label noise, as observed in our experiments, suggests that training data quality significantly impacts posterior accuracy. While this could be partially mitigated with a higher entropy regularization weight, more adaptive joint loss formulations should be explored and specifically adaptive updates of the supervised loss weight $\lambda$. Additionally, active learning or uncertainty-aware sampling strategies could be investigated to guide calibration set selection under a fixed budget. Finally, the training process of our proposed method still involves two separate stages: the joint OT-supervised training and the subsequent amortization of the induced posterior mixture. A more holistic alternative could involve learning the transport mapping directly through a neural OT framework, potentially integrating the matching and inference stages into a single unified process.
Appendix A Datasets clarification
Table 1: Summary of datasets used in the pipeline.
Dataset Observations $\theta$ Labels Usage
$D_{\text{SBI}}$ simulated (${\bm{x}}_{s}$) yes train NPE, NSE
$D_{\text{OT}}$ simulated (${\bm{x}}_{s}$) no train over the OT objective in 3.1, amortization of posterior mixture in 3.2
$D_{\text{calib}}$ real (${\bm{x}}_{r}$) yes train 3.1, MSE objective (and OT)
$D_{u}$ real (${\bm{x}}_{r}$) no train 3.1, OT objective, amortization of posterior mixture in 3.2
Appendix B Semi-balanced Optimal transport solvers.